OpenStudy (anonymous):

let G(x)=5/6 x - 4/5 and find the value of x that corresponds to G(x)=3/4

OpenStudy (anonymous):

we need to find the value of x for which 5/6 x - 4/5=3/4

OpenStudy (anonymous):

given G(x) = 3/4 so solve the eqn 3/4 = 5/6x - 4/5

OpenStudy (anonymous):

solve this eqn.....

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

ok...:)